OpenStudy (magbak):

Rewrite y = x2 + 14x + 29 in general form. (1 point) Rewrite y = 3x2 - 24x + 10 in general form. (1 point) Solve for x: (x - 9)2 = 1 (1 point) Solve for x: x2 + 24x + 90 = 0 (1 point) Solve for x: 2x2 - 4x - 14 = 0 (1 point)

OpenStudy (anonymous):

http://www.purplemath.com/modules/parabola.htm look torward the bottom

OpenStudy (magbak):

I need the explanation please.

OpenStudy (ja1):

Guys we help on this site, we do not give answers nor do we do your homework for you.

OpenStudy (ja1):

http://openstudy.com/code-of-conduct

OpenStudy (magbak):

Ok well you mind your own business please because as you can see I asked for a explanation you illiterate bum.

OpenStudy (ja1):

I was referring to the fact that you posted more than one question, besides you didn't do anything against the CoC @elizabethgamez did.

OpenStudy (magbak):

Was is that I don't understand what you mean.

OpenStudy (ja1):

Yes, but you should post one questions per post but other than that you didn't do anything wrong :)

OpenStudy (magbak):

Oh I apologize for my accusation dear sir.

OpenStudy (ja1):

It's fine I don't mind :)

OpenStudy (magbak):

Thank you.

OpenStudy (anonymous):

\[y = ax ^{2} + bx + c\]\[y = a \left( x ^{2} + \frac{ bx }{ a } \right) + c\]\[y = a \left( x ^{2} + \frac{ bx }{ a } + \frac{ b ^{2} }{ 4a ^{2} } \right) + c - \frac{ b ^{2} }{ 4a }\]\[y = a \left( x + \frac{ b }{ 2a } \right)^{2} + c - \frac{ b ^{2} }{ 4a }\]

OpenStudy (magbak):

Ok can you help me solve one or two of them or at least check two and solve one

OpenStudy (anonymous):

ok. To start, you identify your "a", "b", and "c" and substitute into the last line of the formula. Try one and I'll check it.

OpenStudy (magbak):

1. is y= 3x^2 - 24x + 10 2. Is y equal 3(x-4)^2-38 and solve number 5 please

OpenStudy (magbak):

I mean explain it to me. :)

OpenStudy (anonymous):

ok. Let's take a look at: y= 3x^2 - 24x + 10 Here, this is in "standard form" of: y = ax^2 + bx + c So, a = 3, b = -24, and c = 10 So, substituting into my final line: \[y = 3\left( x - \frac{ 24 }{ (2)(3) } \right)^{2} + 10 - \frac{ (-24)^{2} }{ (4)(3) }\]and now just simplify.

OpenStudy (magbak):

but was I write or no

OpenStudy (magbak):

So but bear with me

OpenStudy (anonymous):

Yes, you got that right. Good job!

OpenStudy (magbak):

and the second one

OpenStudy (magbak):

is it standard form or general form

OpenStudy (anonymous):

Question #5 isn't about general or standard form. It's just about solving for what "x" values will give "0" on the right side. It's a standard form where "y" is already set to "0". 2x^2 - 4x - 14 = 0 -> 2(x^2 - 2x - 7) = 0 and now use the quadratic formula:\[x = \frac{ -b \pm \sqrt{b ^{2} - 4ac} }{ 2a }\]

OpenStudy (magbak):

ohh but 1 and 2 is general form correct

OpenStudy (anonymous):

The form in the problem statement for #1 and #2 are in standard form, but they are asking for those to be re-written in general form, yes. As for #5, you just plug those a, b, and c into the quadratic formula and simplify. a = 1, b = -2, c = -7

OpenStudy (magbak):

So the answers for 1 and 2 are correct yes or no ;) and that will be all. Thank you again for your concern.

OpenStudy (anonymous):

#2 you and I already did and I said that was right. As for #1: y = (x + 7)^2 - 20

OpenStudy (magbak):

ok I appreciate it soooooooooooooooooooooooooooooooooooooooo much Thank youuuuuuuuuuuu I am great full. I hope you have a nice day and again thank you. Bye

OpenStudy (anonymous):

u2! Really nice working with you! @magbak